OpenStudy (jennychan12):

If the substitution u = sqrt(x-1) is made, then the integral from 2 to 5 [(sqrt(x-1)/x]dx = ?

OpenStudy (jennychan12):

\[u = \sqrt{x-1}\] \[\int\limits_{2}^{5} \frac{ \sqrt{x-1} }{ x }dx\]

OpenStudy (jennychan12):

\[du = \frac{ 1 }{ 2\sqrt{x-1} }dx\] => \[dx = 2\sqrt{x-1} du \] is that right?

OpenStudy (jennychan12):

but how would i put that x in the denominator in terms of u?

OpenStudy (jennychan12):

would it be x = u^2 +1 ? then plug into the denominator? oh btw, don't evaluate. just set up. -_- wow, i just kinda answered my own question. how sad

terenzreignz (terenzreignz):

You could figure it out faster than I could even begin to understand how to do it :D

terenzreignz (terenzreignz):

You have to change the limits of integration, though.

OpenStudy (jennychan12):

sorry. yeah i keep getting \[\int\limits_{2}^{5} \frac{ 2u }{ u^2 +1 }du \] can you explain how or why the limits change?

terenzreignz (terenzreignz):

Well, seeing as you're no longer integrating with respect to x, and it WAS x that initially ran from (in this case) 2 to 5, this time it's u that runs.... from u(2) to u(5). That is to say, u, evaluated at x = 2, to u, evaluated at x = 5.

OpenStudy (loser66):

when you set x =...., x =1,---> u =... you must change the limit from x to u. because the original one respect to x , after changing, the function respect to u

OpenStudy (jennychan12):

but why would you evaluate it at u = 2 and u = 5? i'm not really understanding that part :(

terenzreignz (terenzreignz):

Here, let's illustrate the chain rule, and here, F'(x) = f(x) \[\Large \int\limits_{a}^{b}F\circ u(x)u'(x)dx\]

OpenStudy (jennychan12):

yeah i understand that.

terenzreignz (terenzreignz):

Now, we let u=u(x) du=u'(x)dx \[\Large \int\limits_{a}^{b}F(u)du\]

OpenStudy (jennychan12):

ohhh.... that makes sense thanks :)

terenzreignz (terenzreignz):

Aww, I made a mistake in representing... no matter. Suppose F is the derivative of some function G. Then this would be \[\Large G(u)]_{x=a}^{x=b}=G(u(a))-G(u(b))=\int\limits_{u(a)}^{u(b)}F(u)du\]

terenzreignz (terenzreignz):

That's done :)

OpenStudy (jennychan12):

ok, thanks again :)

OpenStudy (zarkon):

did you figure out what your new integral will look like

OpenStudy (jennychan12):

\[\int\limits_{1}^{2} \frac{ 2u^2 }{ u^2+1 }du \] i got the answer, but it took me a little while to figure out why. :/

OpenStudy (zarkon):

good

terenzreignz (terenzreignz):

Well-played :)

OpenStudy (jennychan12):

GG :) sorry again. i understand what you did but why is it G(u(a)) - G(u(b)) ? I thought it's supposed to be G(u(b)) - G(u(a)) ? by the Fundamental Theroem of Calculus ?

terenzreignz (terenzreignz):

Yeah. Typo. Good thing you caught that :)\[\Large G(u)]_{x=a}^{x=b}=G(u(b))-G(u(a))=\int\limits_{u(a)}^{u(b)}F(u)du\] I'd give you another medal, but I can't :D Brilliant error spotting ^.^

OpenStudy (jennychan12):

okay, whew. :o aww:3 haha, thanks.